# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### v4

@@ Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of this result. The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2]. The second proof uses improper integrals of formal power series and is due to this author. We explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, this proof avoids uncountable sets. It is due to this author in [7].
@@
Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity. We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ so that two conditions hold.
@@
the set $x_{n}$ is at most countable (finite or countable). As for uses of mathematical logic in analysis, we quote a referee of [7]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$”. The reader can check [8] for $\mathrm{ACA}_{0}$. Our more ambitious and demanding goal is to obtain an AB proof of the transcendence of $\mathrm{e}$. In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
There are at least three ways to such proof. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
$f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
where $\mathbb{Q}$ is the set of fractions (if real numbers are Dedekind cuts, every such function $f$ is an HMC set), and to develop a sufficient fragment of real analysis for such functions. We realized this way in [7]. Is there any difference in the treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic”. An obvious meta-argument refutes the first claim: if it were the case that “everything works out in the usual way”, then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [7] for details. The second claim is refuted by the fact that in [7] we provide an AB proof. We do not use any technical notion of mathematical logic, and our approach in [7] is therefore very different from the countable treatments of real analysis offered by mathematical logic. For examples of the latter approaches, see Weaver [10, 11].
This analytical digression places the algebraic results of the present article in context. These results are two AB proofs of the transcendence of $\mathrm{e}$. Both are based on formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it already in early versions of [7], and we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC. We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$. It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away. We allow such uses of non-HMC sets in AB proofs for the sake of natural and easily legible formulations.
@@
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. We may work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$. For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
@@
which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we may form the linear combination
@@
In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient
$x^{n}](f\cdot g),\ n\in\mathbb{N}_{0}\,$,
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations. We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any functional values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for any $r$ with $0\leq r\leq k-1$ we indeed have, as $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
@@
3 The proof by formal integration
@@
We next show that shifts behave in desirable ways.
@@
$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
$=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
@@
We will use this identity only with a=1a=-1.
@@
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. We prove three transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
@@
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in Proposition 3.1. $\Box$
@@
We prove the same three integral transformations as before.
@@
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ exists. But then
@@
With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
@@
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
@@
Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
@@
The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and use of many values of formal power series. It is clearly an AB proof.

### v5

@@ Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of this result. The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2]. The second proof uses improper integrals of formal power series and is due to this author. We explain what “semiformal” means.
@@
Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity. We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ in a way satisfying two conditions.
@@
the set $x_{n}$ is at most countable (finite or countable). In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
There are at least four ways to an AB proof of the transcendence of $\mathrm{e}$. For the first two see [7]. The other two ways are the topic of the present article. They are two AB proofs of the transcendence of $\mathrm{e}$ based on formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it already in early versions of [7], and we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC. We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$. It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away. We allow such mentions of non-HMC sets in AB proofs for the sake of natural and legible formulations.
@@
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. We work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$. For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
@@
which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we form the linear combination
@@
In other words, $f(x)\in K\{x\}$ iff $f(x)$ has infinite radius of convergence.
The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient in the product $f\cdot g$,
$[x^{n}](f\cdot g)=\sum_{j=0}^{n}[x^{j}]f\cdot[x^{n-j}]g\,$,}
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations. We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any values of $f(x)$ ($c\in K$)
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}\,.$}
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
@@
3 The proof by semiformal integration
@@
We show that shifts behave in desirable ways.
@@
$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
@@
We use this identity only for a=1a=-1.
@@
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. We obtain three transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
@@
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in item 2 of Proposition 3.1. $\Box$
@@
We obtain for the improper integral the same three integral transformations as before.
@@
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ ($\in\mathbb{R}$) exists. But then
@@
In the other way we argue similarly. $\Box$
@@
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}^{\prime}}$
@@
Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the above properties of $B_{r}$.
@@
The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and the use of many values of formal power series. Clearly it is an AB proof.